An algebra teacher drove by a farmyard full of chickens and pigs. The teacher happened to notice that there were a total of 100 heads and 270 legs. How many chickens were there? how many pigs were there?
Answer by robertb(5830)

c +p =100,
2c + 4p = 270. The top equation gives c = 100-p. Substituting into the bottom equation, 2(100-p) + 4p = 270.
200-2p+4p = 270,
2p = 70,
p = 35, so there are 35 pigs, and 100-35 = 65 chickens.